# Quantum Signatures of Chaos from Free Probability

### v1

@@ 3.2 Free additive convolution
This operation, introduced in voiculescu1986addition , provides a non-commutative analogue of classical convolution in probability theory. The procedure for computing $\rho_{a_{1}+a_{2}}$ is outlined below (see speicher2019lecturenotesfreeprobability ; nica2006lectures ; Potters for comprehensive reviews on the subject, and Zee:1996qu ; zinn1999adding for discussions related to random matrices).
@@
$G(z)=\int_{\mathbb{I}}\textrm{d}\lambda\,\frac{\rho(\lambda)}{z-\lambda}~{},~{}~{}~{}z\in\mathbb{C}\mathbb{I}~{},$ (22)
where $\mathbb{I}$ is the support of the distribution $\rho(\lambda)$ on the real line. For example, for a $\mathcal{D}\times{\mathcal{D}}$ random matrix Hamiltonian ($H$), consider its resolvent defined as, $G_{N}(z)=\frac{1}{{\mathcal{D}}}{\rm Tr}\,\frac{1}{z-H}=\frac{1}{{\mathcal{D}}}\sum_{i}\frac{1}{z-\lambda_{i}}$. Now averaging over the random matrix ensemble, and taking the limit ${\mathcal{D}}\rightarrow\infty$, we have the Cauchy transform of the eigenvalue density: $\mathbb{E}(G_{N}(z))\equiv G(z)=\int_{\mathbb{I}}\textrm{d}\lambda\,\frac{\rho(\lambda)}{z-\lambda}$.  [^8] For a self-adjoint operator, say $\mathcal{O}$, the Cauchy transformation of its eigenvalue distribution can also be thought of as the moment generating function of this operator, i.e., $G_{\mathcal{O}}(z)=\sum_{k=0}^{\infty}\frac{\varphi(\mathcal{O}^{k})}{z^{k+1}}$, where $\varphi(\mathcal{O}^{k})$ denotes the $k$-th order moment of the operator (see Appendix A). E.g., for a random matrix Hamiltonian, we have, $\varphi(H^{k})=\mathbb{E}({\rm Tr}\,(H^{k}))=\int\rho(\lambda)\lambda^{k}\textrm{d}\lambda$. The expansion of the $G(z)$ indicates that it is well-behaved at $\infty$, and therefore, having the knowledge of this function near infinity is equivalent to having the knowledge of all the moments of the operator.
@@
$G^{-1}(z)\equiv\mathcal{B}(z)=\frac{1}{z}+R(z),$ (24)
@@
Since the Cauchy transform, and consequently the eigenvalue density, may admit multiple solutions, the physically relevant solution must be chosen by ensuring positivity within the appropriate domain. There can also be solution for $G_{\rho_{a_{1}}\boxplus\rho_{a_{2}}}(z)$ which are complex conjugates of each other. In that case, one can take the absolute value of any one of these conjugate solutions to get the correct eigenvalue density (see the example of convolution of eigenvalue density of the sum of two free spin-1 operators that we discuss below).
@@
$\rho_{\sigma}(\lambda)=\frac{\delta(\lambda-1)+\delta(\lambda+1)}{2}.$ (28)
Since both $\rho_{\sigma_{1}}$ and $\rho_{\sigma_{2}}$  [^10] To simplify the notations, below, we denote $\rho_{\sigma_{1}}$ and $\rho_{\sigma_{2}}$ as $\rho_{1}$ and $\rho_{2}$, respectively. The free additive convolution of them is denoted as $\rho_{3}$, i.e., $\rho_{3}=\rho_{1}\boxplus\rho_{2}$. have compact supports on $\mathbb{R}$, one can obtain the Cauchy transformation (22) of this distribution to be
$G_{\rho_{1}}(z)=\frac{z}{z^{2}-1}.$ (29)
@@
$\rho_{\Sigma}(\lambda)=\frac{\delta(\lambda-1)+\delta(\lambda)+\delta(\lambda+1)}{3}.$ (34)
@@
From this, we first obtain $\mathcal{B}_{\rho_{3}}(z)=R_{\rho_{3}}(z)+\frac{1}{z}$ and calculate the corresponding Cauchy transformation by using the inversion formula $\mathcal{B}_{\rho_{3}}(G_{\rho_{3}}(z))=z$. There are multiple solutions of this equation for the Cauchy transform $G_{\rho_{3}}(z)$. To obtain the correct eigenvalue density, we first plot the function $\mathcal{B}_{\rho_{3}}(G_{\rho_{3}}(z))(=z)$, shown in Fig. 4. There is a region on the vertical axis (from around $-1.92$ to $1.92$) where the function $\mathcal{B}_{\rho_{3}}$ does not attain real values, and hence, this corresponds to the region where the eigenvalues of the sum operator lie. Furthermore, as we have shown in the inset of Fig. 4, this exists between the local minimum and maximum of the function $\mathcal{B}_{\rho_{3}}(G_{\rho_{3}})$. More precisely, we determine the approximate edges of the spectrum of $\Sigma_{3}$ to be between $-1.9227$ to $1.9227$.
@@
Having determined the range of the spectrum, we now determine the density of eigenvalues of the convoluted operator. As we are eventually looking for a convoluted probability distribution, it should be positive definite in an interval in $\mathbb{R}$, which, in this case, is approximately between $-1.9227$ to $1.9227$.
@@
This is plotted in Fig. 5. Furthermore, as one can check, here taking the limit $\epsilon\rightarrow 0$ gives the same distribution as directly substituting $\epsilon=0$, it is possible to obtain the following analytical expression for the distribution shown in Fig. 5 as well,
@@
Figure 5: Convolved distribution for the eigenvalue density of the sum of two free spin-1 operators according to free probability theory. The approximate edges of the spectrum, as determined from the plot of the function $\mathcal{B}_{\rho_{3}}(G_{\rho_{3}}(z))(=z)$ in Fig. 4 are, respectively, -1.9227 and 1.9227.

### v3

@@ 3.1 Statistical Independence in Quantum Mechanics: Freeness
3.2 Asymptotic freeness and the vanishing of OTOCs
Under these conditions, we expect that for two operators $A,B\in\mathcal{A}$, if the dynamics is chaotic, the time-evolved operator $B(t)$ becomes asymptotically free from $A(0)$ with respect to the state (22) at sufficiently large times. This expectation is motivated by the fact that in a chaotic system, the time evolution operator $e^{-iHt}$ can be approximated by a Haar-random unitary, i.e., [^9] Note that when the quantum dynamics modeled by a random unitary matrix from the Haar measure of the unitary groups, since there is no notion of time-translation symmetry, energy is not conserved in such a dynamics. A more physical approach, which is closer to the dynamics generated by a generic chaotic quantum system with energy conservation, would be to consider unitary evolution $U=e^{-itH}$ generated random matrix Hamiltonian $H$ drawn from some ensemble of (time-independent) Hamiltonians. The time evolution operators form an ensemble $\mathcal{E}(t)$, which, under time evolution, would gradually evolve into a complicated unitary ensemble, though not exactly Haar random, in the long-time limit. In Cotler:2017jue the time evolution by an ensemble of GUE Hamiltonians was considered, and its approach to the Haar random unitary was quantified by examining when an approximate $k$-design is formed by computing the so-called frame potentials. Interestingly, it was found that time evolution by the GUE Hamiltonian does form an approximate $k$-design at an intermediate time scale; however, subsequently it then deviates from being an approximate $k$-design at later times.
$e^{-iHt}\approx U_{\text{Haar}}.$
3.3 Incorporating Freeness into the Quantum Ergodic Hierarchy
F-systems.
A quantum dynamical system $(\mathcal{A},\varphi,T_{t})$ is said to be an F-system or to possess the F-property if, for any operators $A,B\in\mathcal{A}$, the operators $A$ and $B(t)$ are free in the limit $t\to\infty$. This implies the vanishing of all mixed free cumulants involving $A$ and $B(t)$:
$\lim_{t\rightarrow\infty}\kappa_{n}(A,B(t),A,B(t),\ldots,A,B(t))=0\,.$ (23)
We also refer to the F-property as pairwise freeness.
Note that the F-property implies quantum 2-mixing, but it does not imply quantum $n$-mixing. Consider $n$ Heisenberg operators $A_{1}(t_{1}),A_{2}(t_{2}),\ldots,A_{n}(t_{n})$, and take the limit $t_{i+1}-t_{i}\to\infty$ for all $i=1,\ldots,n-1$. The F-property implies that, for any $i$, $A_{i+1}(t_{i+1})$ is free from $A_{i}(t_{i})$ as $t_{i+1}-t_{i}\to\infty$, which in turn implies that $A_{1}(t_{1}),A_{2}(t_{2}),\ldots,A_{n}(t_{n})$ are pairwise free. Pairwise freeness, however, does not imply mutual freeness, which is necessary for quantum $n$-mixing (16). It is interesting to compare the F-property with the notion of a quantum K-system, which implies quantum $n$-mixing (see Theorem 2.3 in Narnhofer:2001 ), but does not imply the vanishing of out-of-time-order correlators (see Example 2.2 in Narnhofer:2001 ). In particular, this shows that quantum K-property, as a mathematical condition, is not enough for the emergence of freeness. Therefore, while we expect that physical examples of quantum K-systems may also display the F-property, and vice versa, the mathematical definitions of these properties are not equivalent; neither condition is strictly stronger or weaker than the other.
In Figure 4, we illustrate the quantum and classical ergodic hierarchies and their consequences for quantum systems with finite dimensional Hilbert spaces.
3.4 Asymptotic freeness produced by time evolution
In this section, we use the formulas provided in Appendix D to gain intuition about the time scale when the mixed free cumulants of two deterministic operators vanish under time evolution, so that the operators approach freeness [^11] In the context of Anosov flows, the exponential decay of correlation functions during the approach to equilibrium is governed by the so-called Pollicott–Ruelle resonances (see 2017pollicottruelleresonances for a review). One also expects these resonances to control the approach to equilibrium of OTOCs  Polchinski:2015cea . Since freeness emerges when correlation functions and OTOCs involving centered operators vanish, we expect the time scale for the onset of freeness to also be governed by the Pollicott–Ruelle resonances in systems in which such concepts are well-defined. We thank Zhenbin Yang for discussions on this point..
To start with, first we consider the case of the time evolution operator modeled by a random unitary matrix that is drawn from the Haar measure on the unitary group. It is known that the operators $A$ and $B_{U}=U^{\dagger}BU$ are free when $U$ is a unitary matrix belonging to the Haar measure of the unitary group (where $A$ and $B$ are two deterministic ${\mathcal{D}}\times{\mathcal{D}}$ matrices which have convergent densities in the limit $\mathcal{D}$$\rightarrow\infty$). To see this from the explicit calculation of the moments, first consider the two-point function between the operators $A$ and $B_{U}=U^{\dagger}BU$, which can be written as
$\big{<}AB_{U}\big{>}=\langle A\rangle\langle B_{U}\rangle~.$ (24)
This expression is the same one would expect when the operators $A$ and $B_{U}$ are free. Next, considering the $4$-point OTOC one can show that Roberts:2016hpo ; Yoshida:2017non
$\langle AB_{U}AB_{U}\rangle=\frac{{\mathcal{D}}^{2}}{{\mathcal{D}}^{2}-1}\Bigg{[}\langle A^{2}\rangle\langle B\rangle^{2}+\langle A\rangle^{2}\langle B^{2}\rangle-\langle A\rangle^{2}\langle B\rangle^{2}-\frac{1}{{\mathcal{D}}^{2}}\langle A^{2}\rangle\langle B^{2}\rangle\Bigg{]}~.$ (25)
Clearly, in the limit $\mathcal{D}$$\rightarrow\infty$, the final term in the above expression vanishes, and it reduces to the one we get when the operators $A$ and $B_{U}$ are free. One can follow a similar procedure to show that the higher-order correlators also reduce to the ones predicted by the free probability theory in the $\mathcal{D}$$\rightarrow\infty$ limit.
To understand the effect of time-evolution by a generic Hamiltonian on two free variables, let us consider the following $2$-point correlation function between two deterministic matrices $A$ and $B$, $\langle Ae^{iHt}Be^{-iHt}\rangle$, where $H$ is a generic Hamiltonian, not necessarily belonging to some random matrix ensemble. Now, consider rotating one of the two operators by a random rotation by a unitary matrix $U$ belonging to the Haar measure of the unitary group, and compute the average of the correlation function over the group. A simple calculation shows that,
$\int\text{d}U~\langle Ae^{iHt}B_{U}e^{-iHt}\rangle=\langle A\rangle\langle B_{U}\rangle~,$ (26)
i.e., the average two-point correlation factorises, and is independent of time, irrespective of the Hamiltonian. Similarly, computing the following four-point function,
$\int\text{d}U~\langle Ae^{iHt}B_{U}e^{-iHt}Ae^{iHt}B_{U}e^{-iHt}\rangle=\frac{{\mathcal{D}}^{2}}{{\mathcal{D}}^{2}-1}\Bigg{[}\langle A^{2}\rangle\langle B\rangle^{2}+\langle A\rangle^{2}\langle B^{2}\rangle-\langle A\rangle^{2}\langle B\rangle^{2}-\frac{1}{{\mathcal{D}}^{2}}\langle A^{2}\rangle\langle B^{2}\rangle\Bigg{]}~,$ (27)
we again see that it becomes time independent, and in the large $\mathcal{D}$} limit, $\mathcal{D}$$\rightarrow\infty$ matches with the free probability expression one gets when the variables $A$ and $B_{U}$ are free. Therefore, time evolution by a generic Hamiltonian does not change the mutual freeness of two initially free variables. This suggests that chaos should be understood as a mechanism that drives initially non-free operators toward asymptotic freeness, whereas initially free variables remain free under time evolution.
Time evolution generated by random matrices and the spectral form factor.
Next, consider the quantum dynamics generated by a random matrix, and the correlation function between the operators $A$ and $B(t)=e^{iHt}Be^{-iHt}$. As we have discussed, for a random matrix Hamiltonian drawn from the GUE, the two-point function can be written in terms of the spectral form factor as (see eq. (115))
$\begin{split}\langle AB(t)\rangle_{\text{GUE}}=\int\text{d}U~\langle A_{U}e^{iHt}B_{U}e^{-iHt}\rangle_{\text{GUE}}=\langle AB\rangle\frac{\mathcal{R}_{2}(t)-1}{{\mathcal{D}}^{2}-1}+\langle A\rangle\langle B(t)\rangle\frac{{\mathcal{D}}^{2}-\mathcal{R}_{2}(t)}{{\mathcal{D}}^{2}-1}~.\end{split}$ (28)
We want to see from this formula and the known time-evolution pattern of the spectral form factor when the above correlation function goes to $\langle A\rangle\langle B(t)\rangle$ consistent with the free probability prediction. First, note that, as expected, at the start of the evolution at $t=0$, when $\mathcal{R}_{2}(t)={\mathcal{D}}^{2}$, the second term above vanishes, and the first one is equal to $\langle AB\rangle$. For $t>0$, which of the two terms dominates depends on the ratio, $r_{f}=(\mathcal{R}_{2}(t)-1)/({\mathcal{D}}^{2}-\mathcal{R}_{2}(t))$. For instance, before reaching the dip, when $\mathcal{R}_{2}(t)$ crosses a value of $\mathcal{D}$}, the above ratio is $\mathcal{O}({\mathcal{D}}^{-1})$ so that the second term above dominates in the limit $\mathcal{D}$$\rightarrow\infty$ and the two-point correlation function matches approximately with the free probability prediction. Furthermore, at dip time, $t_{d}\approx\sqrt{{\mathcal{D}}}$, and $\mathcal{R}_{2}(t)\approx\sqrt{{\mathcal{D}}}$ Cotler:2017jue , so that the ratio goes to zero faster ($\mathcal{O}({\mathcal{D}}^{-3/2})$) in the limit $\mathcal{D}$$\rightarrow\infty$. However, for times larger than $t>t_{d}$, when the spectral form factor shows a ramp, the decay is slower, even though it still goes to zero when the limit $\mathcal{D}$$\rightarrow\infty$ is considered. Finally, when the spectral form factor reaches the plateau, at time $t_{P}>2{\mathcal{D}}$, $\mathcal{R}_{2}(t)$ reaches a constant value $\mathcal{O}({\mathcal{D}})$, and the ratio $r_{f}$ vanishes as ${\mathcal{D}}^{-1}$ in the large $\mathcal{D}$} limit for all subsequent times. To summarise, for time evolution generated by the a random matrix Hamiltonian drawn from GUE, for times scales larger than at least the time when the spectral form factor takes a value $\mathcal{O}({\mathcal{D}})$, the $2$-point correlation function of two deterministic matrices factorises and matches with the free probability prediction in the large $\mathcal{D}$} limit.
One can, in principle, use a similar procedure to the one reviewed in the Appendix. D to obtain the four-point OTOC by employing the formula for the integral over the Haar random unitary group. However, due to the fact that this integral gives a large number of terms, it is inconvenient to write the exact expression of the OTOC. Rather than following this approach, in the following, we employ techniques from free probability theory to compute the behavior of the OTOC valid in the large $\mathcal{D}$ limit.
OTOCs for random matrices from free probability.
Consider the OTOC $\langle A(0)B(t)A(0)B(t)\rangle$ between two operators $A$ and $B(t)$, with the time evolution being generated by a random matrix Hamiltonian of rank $\mathcal{D}$} drawn from GUE, in the limit $\mathcal{D}$$\rightarrow\infty$. We shall use some results from free probability theory to obtain an expression for this correlator. [^12] For previous discussions in the literature on the OTOCs for time evolution generated by random matrices, see Cotler:2017jue ; you2018entanglement ; Vijay:2018qmx ; Bellitti:2019hhn . For simplicity, we consider the case that the operators $A$ and $B$ are initially non-overlapping, i.e., $\langle AB\rangle=\langle A\rangle\langle B\rangle$ [^13] E.g., for the case of spin systems, these could be the Pauli spin operators at different sites., and they have vanishing first order moment, i.e., $\langle A\rangle=0=\langle B\rangle$. Expanding the OTOC, we have
$\text{OTOC}(t)=\sum_{k,l,m,n}\frac{(it)^{k+m}(-it)^{l+n}}{k!~l!~m!~n!}\langle AH^{k}BH^{l}AH^{m}BH^{n}\rangle~.$ (29)
Now, to compute the mixed moment $m=\langle AH^{k}BH^{l}AH^{m}BH^{n}\rangle$, we use the fact the large random matrix $H$ of rank $\mathcal{D}$} is free from both the deterministic matrices $A$ and $B$ of the same rank, in the limit $\mathcal{D}$$\rightarrow\infty$ mingo2017free . Therefore, considering the expansion of the moment in terms of the free cumulants, we see that the only non-zero contribution comes from the diagram in Fig. 3, so that it can be written as
$m=\kappa_{4}(ABAB)\langle H^{k}\rangle\langle H^{l}\rangle\langle H^{m}\rangle\langle H^{n}\rangle~.$ (30)
Now using the relation $\langle ABAB\rangle=\kappa_{4}(ABAB)+2\kappa_{2}(AB)$ valid for two operators $A$ and $B$ having vanishing first moments, along with the fact that for a Hamiltonian drawn from the GUE, we have [^14] Since, $H$ is a Gaussian random matrix, $\langle H^{n}\rangle$ is non-zero only when $n$ is even. $\langle H^{n}\rangle=C_{n/2}$, $C_{n}$ being the Catalan numbers $C_{n}=\frac{(2n)!}{n!(n+1)!}$, we have
$\text{OTOC}(t)=\langle ABAB\rangle\sum_{k,l,m,n}\frac{(it)^{2(k+m)}(-it)^{2(l+n)}}{(2k)!~(2l)!~(2m)!~(2n)!}C_{k}C_{l}C_{m}C_{n}~=\langle ABAB\rangle\bigg{(}\frac{I_{1}(2it)}{it}\bigg{)}^{4}~.$ (31)
Therefore, the OTOC shows power law decay at early times, and eventually vanishes at later times, consistent with the free probability prediction when the operators $A$ and $B(t)$ are free. We also emphasize that even though they match with the free probability prediction at early times, the numerical result for finite-dimensional random matrices, as well as the result obtained using the formula for the higher-order averages over the Haar random ensemble, do not go to zero at later times, rather saturate to a small value. Only in the large $\mathcal{D}$} limit does the numerical prediction for the OTOC vanish at late times and match with the free probability prediction.
Figure 3: The diagram that contributes to the moment $m=\langle AH^{k}BH^{l}AH^{m}BH^{n}\rangle$ for two traceless operators $A$ and $B$ which have zero initial overlap. Since the large Gaussian matrices $H$ is free from both the deterministic matrices $A$ and $B$, all their mixed free cumulant vanish, and hence the contribution from this diagram is $m=\kappa_{4}(ABAB)\langle H^{k}\rangle\langle H^{l}\rangle\langle H^{m}\rangle\langle H^{n}\rangle$.
Before concluding this section, we discuss the following point regarding the difference between the free probability prediction for the OTOC and the one that exists in the literature, computed e.g., using the higher-order moments of the Haar random ensemble. This will highlight the importance of the limit $\mathcal{D}$$\rightarrow\infty$ as well as helps us to understand the role played by the eigenvalue correlation in the random matrix Hamiltonian, which generates the time evolution. First, we compare the following leading order expression for the OTOC obtained in Cotler:2017jue
$\text{OTOC}(t)\simeq\langle ABAB\rangle\frac{\mathcal{R}_{4}(t)}{{\mathcal{D}}^{4}}~,~\mathcal{R}_{4}(t)=\langle|\text{Tr}(e^{-iHt})|^{4}\rangle=\int\mathcal{D}\lambda\sum_{k,l,m,n}e^{i(E_{k}+E_{l}-E_{m}-E_{n})t}~,$ (32)
with the OTOC obtained in (31). Here $\mathcal{R}_{4}(t)$ is the $4$-point spectral form factor at infinite temperature, its expression obtained using the so-called box-approximation can be written as Cotler:2017jue
$\frac{\mathcal{R}_{4}(t)}{{\mathcal{D}}^{4}}=r_{1}^{4}(t)+2{\mathcal{D}}^{-2}r_{2}^{2}(t)-4r_{2}(t)\big{(}{\mathcal{D}}^{-2}-{\mathcal{D}}^{-3}\big{)}-7{\mathcal{D}}^{-3}r_{2}(2t)+4{\mathcal{D}}^{-3}r_{2}(3t)+(2{\mathcal{D}}^{-2}-{\mathcal{D}}^{-3})~,$ (33)
where $r_{1}(t)$ is the Fourier transform of the density of states of the Hamiltonian, while $r_{2}(t)$ is the double Fourier transform of the connected part of the two-point correlation function (which is the sine kernel for GUE mehta1991random ; Cotler:2016fpe ), and their expressions are given by,
$r_{1}^{2}(t)=|\langle e^{-itH}\rangle|^{2}=\bigg{(}\frac{J_{1}(2t)}{t}\bigg{)}^{2}~,~~\text{and}~~r_{2}(t)=\begin{cases}1-\frac{t}{2{\mathcal{D}}}~,~&~\text{for}~~t<2{\mathcal{D}}\\ 0~,~&~\text{for}~~t>2{\mathcal{D}}~.\end{cases}$ (34)
Thus, $r^{2}_{1}(t)$ is the part of the spectral form factor that comes from the disconnected part of the two-point correlation function. Now we can observe from the expression in Eq. (33) that in the limit $\mathcal{D}$$\rightarrow\infty$, the dominant contribution to the OTOC in eq. (32) comes from the term $r_{1}^{4}(t)=|\langle e^{-itH}\rangle|^{4}$, i.e. the disconnected part of the spectral form factor, and hence matches with the expression for the OTOC obtained from the free probability theory in eq. (31). In other words, the correlation between the eigenvalues of a random matrix Hamiltonian does not seem to have an affect on the OTOC in the limit when the rank of the Hamiltonian goes to infinity. [^15] One can easily check that this is also true for two-point functions. In the free probability computation in eq. (31) (and the Fig. 3), this is a direct consequence of the fact that crossing partitions do not contribute to the moment expansion.
4 Eigenvalue statistics of operators as a signature of asymptotic freeness
The fact that freeness can only emerge in the large $\mathcal{D}$ [^16] We remind the reader that $\mathcal{D}$ denotes the rank of the operators in the algebra. limit can be understood from an algebraic perspective, particularly in the context of the quantum ergodic hierarchy. First, observe that freeness implies 2-mixing. Since mixing is a property that arises only for type II or III algebras, it follows that freeness can occur only for type II or type III algebras. For finite $\mathcal{D}$, however, the algebra of observables is always of type I, and therefore, freeness cannot occur. As a result, mixed free cumulants of centered variables ($A-\langle A\rangle$ and $B(t)-\langle B(t)\rangle$) will not generically vanish for finite $\mathcal{D}$. However, if the dynamics is chaotic, these cumulants are expected to approach a small residual value that scales inversely with the number of degrees of freedom in the system. This expectation was indeed shown to be true for OTOCs in Huang:2017fng . Furthermore, if two operators $A$ and $B(t)$ are free, the eigenvalue density of their sum $\rho_{A(0)+B(t)}$ is expected to follow the free convolution prediction, which can emerge even for finite, but sufficiently large $\mathcal{D}$, and serves as a “smoking gun” of asymptotic freeness Chen:2024zfj . These concepts are illustrated schematically in Figure 4.
Figure 4: Illustration of the quantum and classical ergodic hierarchies. NN denotes the number of degrees of freedom of the system. Freeness (F-property) is stronger than quantum 2-mixing, but it is neither stronger nor weaker than the quantum K-property. The corresponding von Neumann algebraic structures associated with different degrees of quantum chaos are indicated. Note that the quantum hierarchy is only meaningful in the $1/N\rightarrow 0$ limit, as for any finite $N$, the algebra remains type I, preventing even ergodicity. Nonetheless, a finite-NN system can exhibit signatures of its asymptotic behavior. For example, the emergence of operator spectral statistics governed by free probability theory serves as a robust indicator of asymptotic freeness.
@@ 4.1 Free additive convolution
This operation, introduced in voiculescu1986addition , provides a non-commutative analogue of classical convolution in probability theory. The procedure for computing $\rho_{a_{1}+a_{2}}$ is outlined below (see speicher2019lecture ; nica2006lectures ; Potters ; tao2012topics for comprehensive reviews on the subject, and Zee:1996qu ; zinn1999adding for discussions related to random matrices).
@@
$G(z)=\int_{\mathbb{I}}\textrm{d}\lambda\,\frac{\rho(\lambda)}{z-\lambda}~,~~~z\in\mathbb{C}\,$, (36)
where $\mathbb{I}$ is the support of the distribution $\rho(\lambda)$ on the real line. For example, for a $\mathcal{D}\times{\mathcal{D}}$ random matrix Hamiltonian ($H$), consider its resolvent defined as, $G_{N}(z)=\frac{1}{{\mathcal{D}}}{\rm Tr}\,\frac{1}{z-H}=\frac{1}{{\mathcal{D}}}\sum_{i}\frac{1}{z-\lambda_{i}}$. Now averaging over the random matrix ensemble, and taking the limit $\mathcal{D}$$\rightarrow\infty$, we have the Cauchy transform of the eigenvalue density: $\mathbb{E}(G_{N}(z))\equiv G(z)=\int_{\mathbb{I}}\textrm{d}\lambda\,\frac{\rho(\lambda)}{z-\lambda}$.  [^17] For a self-adjoint operator, say $\mathcal{O}$, the Cauchy transformation of its eigenvalue distribution can also be thought of as the moment generating function of this operator, i.e., $G_{\mathcal{O}}(z)=\sum_{k=0}^{\infty}\frac{\varphi(\mathcal{O}^{k})}{z^{k+1}}$, where $\varphi(\mathcal{O}^{k})$ denotes the $k$-th order moment of the operator (see Appendix A). E.g., for a random matrix Hamiltonian, we have, $\varphi(H^{k})=\mathbb{E}({\rm Tr}\,(H^{k}))/\mathcal{D}=\int\rho(\lambda)\lambda^{k}\textrm{d}\lambda$. The expansion of the $G(z)$ indicates that it is well-behaved at $\infty$, and therefore, having the knowledge of this function near infinity is equivalent to having the knowledge of all the moments of the operator.
@@
$G^{-1}(z)\equiv\mathcal{B}(z)=\frac{1}{z}+R(z)~,$ (38)
@@
Since the Cauchy transform, and consequently the eigenvalue density, may admit multiple solutions, the physically relevant solution must be chosen by ensuring positivity within the appropriate domain. There can also be solution for $G_{\rho_{a_{1}}\boxplus\rho_{a_{2}}}(z)$ which are complex conjugates of each other. In that case, one can take the absolute value of any one of these conjugate solutions to get the correct eigenvalue density (see the example of the convolution of the eigenvalue density of the sum of two free spin-1 operators that we discuss below).
@@
$\rho_{\sigma}(\lambda)=\frac{\delta(\lambda-1)+\delta(\lambda+1)}{2}~.$ (42)
Since both $\rho_{\sigma_{1}}$ and $\rho_{\sigma_{2}}$  [^19] To simplify the notations, below, we denote $\rho_{\sigma_{1}}$ and $\rho_{\sigma_{2}}$ as $\rho_{1}$ and $\rho_{2}$, respectively. The free additive convolution of them is denoted as $\rho_{3}$, i.e., $\rho_{3}=\rho_{1}\boxplus\rho_{2}$. This convention is followed through out the paper. have compact supports on $\mathbb{R}$, one can obtain the Cauchy transformation (36) of this distribution to be
$G_{\rho_{1}}(z)=\frac{z}{z^{2}-1}~.$ (43)
@@
$\rho_{\Sigma}(\lambda)=\frac{\delta(\lambda-1)+\delta(\lambda)+\delta(\lambda+1)}{3}~.$ (48)
@@
From this, we first obtain $\mathcal{B}_{\rho_{3}}(z)=R_{\rho_{3}}(z)+\frac{1}{z}$ and calculate the corresponding Cauchy transformation by using the inversion formula $\mathcal{B}_{\rho_{3}}(G_{\rho_{3}}(z))=z$. There are multiple solutions of this equation for the Cauchy transform $G_{\rho_{3}}(z)$. To obtain the correct eigenvalue density, we first plot the function $\mathcal{B}_{\rho_{3}}(G_{\rho_{3}}(z))(=z)$, shown in Fig. 6. There is a region on the vertical axis (from around 1.92-1.92 to $1.92$) where the function $\mathcal{B}_{\rho_{3}}$ does not attain real values, and hence, this corresponds to the region where the eigenvalues of the sum operator lie. Furthermore, as we have shown in the inset of Fig. 6, this exists between the local minimum and maximum of the function $\mathcal{B}_{\rho_{3}}(G_{\rho_{3}})$. More precisely, we determine the approximate edges of the spectrum of $\Sigma_{3}$ to be between 1.9227-1.9227 to $1.9227$.
@@
Having determined the range of the spectrum, we now determine the density of eigenvalues of the convolved operator. As we are eventually looking for a convolved probability distribution, it should be positive definite in an interval in $\mathbb{R}$, which, in this case, is approximately between 1.9227-1.9227 to $1.9227$.
@@
This is plotted in Fig. LABEL:fig:_spin1_plot. Furthermore, as one can check, here taking the limit $\epsilon\rightarrow 0$ gives the same distribution as directly substituting $\epsilon=0$, it is possible to obtain the following analytical expression for the distribution shown in Fig. LABEL:fig:_spin1_plot as well,
@@
Figure 7: Convolved distribution for the eigenvalue density of the sum of two free spin-1 operators according to free probability theory. The approximate edges of the spectrum, as determined from the plot of the function $\mathcal{B}_{\rho_{3}}(G_{\rho_{3}}(z))(=z)$ in Fig. 6, are, respectively, -1.9227 and 1.9227.